Understand the price elasticity of demand formula

2. Draw a perfectly elastic and perfectly inelastic demand curve and label each

3. Be able to identify whether demand is elastic or inelastic given changes in quantity and price

4. Be able to calculate percentage change using the midpoint formula and be able to apply it to calculate the price elasticity of demand

5. Know the determinants of the price elasticity of demand and be able to identify how they change price elasticity

6. Understand how to identify whether two goods are complements, substitutes, normal, inferior, necessity or luxury given cross price elasticity or income elasticity

7. How do you calculate total revenue? If an ice cream shop charges $5 for an ice cream cone and they sell 20 cones, what is their total revenue?

8. When is total revenue maximized?

9. What is the definition of marginal utility and how do you calculate it?

10. What is the law of diminishing marginal utility?

11. How do consumers maximize total utility?

12. Be able to identify the optimal consumption bundle given a chart and on a graph?

13. Define behavioral economics

14. Define sunk costs

15. Understand what an indifference curve and budget constraint show and how they change when prices and income change.

16. What is technology?

17. What is the difference between the short run and the long run?

18. What does the production function show?

19. Understand and be able to calculate the marginal product labor.

20. Identify where diminishing marginal returns occurs as related to labor.

21. Be able to draw and label a graph with marginal cost (MC), marginal revenue (MR), average total cost (ATC), average variable cost (AVC) and average fixed cost (AFC).

22. What does an isoquant show?

23. ATC = AFC + ______

24. What are the characteristics of a perfectly competitive market?

25. Understand the concept of price takers related to PC firms

26. What is the relationship between demand, price andmarginal revenue in a PC market?

27. What are the profit maximizing rules for all firms and for PC firms?

28. Be able to read a graph with the above curves and calculate the following: profit maximizing output level, total fixed cost, minimum price to produce at, profit or loss at specific price level.

29. Identify at what price a firms makes a profit, breaks even, experiences a loss using a graph
Answer #1

0 a formula Price elasticity of demand form V. change in Quantity demanded T. change in price q = Quantity earlier O, - Q x 1P, = 2 P2=3 9 = 100 Q = 80 Mid-point Method 80--10 90=(80 +100/2 (Q2+Q,)/2 POP, (P₂ + Pid2 3-2 2.5=(3+2/2 = - 20%, X100 7.5x1